# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide QED

### v1

@@ I Introduction
Recently, there has been a revived interest in superradiance [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], mainly stimulated by the development of new experimental platforms, in which superradiant effects can be explored at an unprecedented level of control, and potentially be exploited in applications such as superradiant lasers [43, 44, 45] or quantum metrology [46, 47]. Of particular relevance in this context are waveguide QED systems [48, 49, 50], where the emitted photons are confined to a one-dimensional (1D) channel and atom-atom correlations can be established efficiently and over long distances [16, 18, 19, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Specifically, when emitters are placed in a lattice with a spacing that matches an exact multiple of the resonant wavelength, all the emitted photons can interfere fully constructively, thereby recovering the features of the original Dicke model in an extended setting. However, in real systems, the positions of the atoms as well as their frequencies fluctuate, giving rise to competing interference patterns once the disorder becomes sufficiently strong. This raises an important fundamental question: Does Dicke superradiance survive in disordered quantum systems, and how does disorder affect its characteristic features?
Figure 1: (a) Sketch of a disordered waveguide QED system, where NN excited two-level atoms are located at random positions $z_{j}=j\times\lambda_{0}+\delta_{j}$ along a 1D photonic channel and undergo collective spontaneous decay. As shown in the inset, the displacements $\delta_{j}$ from exact multiples of the resonant wavelength $\lambda_{0}$ are chosen randomly from the interval $\Theta/2,\Theta/2)$ to interpolate between a regular array ($\Theta=0)$ and a uniformly distributed gas of atoms ($\Theta\gg 1$). (b) Illustration of the two ordering patterns (red and blue), which are generated spontaneously between the atomic dipoles during the superradiant decay. The dipoles self-align according to their distances, with $\phi_{j}\approx\phi_{1}+k_{0}(z_{j}-z_{1})$ (blue arrows) or $\phi_{j}\approx\phi_{1}-k_{0}(z_{j}-z_{1})$ (red arrows), which even for strong disorder yields constructive interference and enhanced collective emission either to the left or to the right.
@@
By investigating the maximal emission rate and the peak emission time for an increasing number of atoms $N$, we demonstrate that even for strong disorder, the Dicke scaling is asymptotically recovered. There are, however, subtle differences in the numerical prefactors and in how this limit is approached. To explain these observations, we introduce a product-state ansatz for the configuration of atomic dipoles at the time of the emission peak, which provides a tight upper bound for the maximal collective decay rate. Remarkably, this bound is not only close to the numerically exact values, but also reveals a simple and intuitive mechanism for robustness: the atomic spins organize spontaneously in random directions, but with their relative orientations being determined by their relative locations along the waveguide [Fig. 1(b)]. This emergent spin ordering, without any external control, enables constructive interference in the decay channels for any atomic configuration, which enables the persistence of the $N^{2}$ scaling regardless of the disorder strength.
@@ II Model
We consider a set of $N$ TLAs, which are located at positions $z_{j}$ along a 1D photonic waveguide [see Fig. 1(a)]. The TLAs have a ground state $|g\rangle$ and an excited state $|e\rangle$, which are separated by a transition frequency $\omega_{0}$. Under the assumption that over a sufficiently broad frequency range around $\omega_{0}$ the waveguide exhibits an approximately linear dispersion relation with group velocity $v_{\rm g}$, we can trace out the waveguide modes and describe the remaining dynamics of the atoms by a master equation of the form (see, for example, [55, 56])
@@
In our study, the offset phases $\xi_{j}=k_{0}\delta_{j}$ are randomly chosen from a uniform distribution in the interval $\Theta/2,\Theta/2)$, such that $\Theta$ quantifies the degree of disorder. Additional simulations for an ensemble with Gaussian-distributed $\xi_{j}$ are performed for comparison, ensuring that the general conclusions do not depend on any specifics of the disorder.
@@
where $J_{\pm}$ are collective jump operators and $\Gamma_{\pm}=\gamma N(1\pm|c|)/2$ are the corresponding decay rates. Here, we have introduced the parameter
@@
which characterizes the overlap between the original left- and right-going decay modes. It assumes a value of $c=1$ in the absence of disorder, i.e. $\xi_{j}=0$, and a value of $c\simeq 0$ when the atoms are homogeneously distributed. The short-time dynamics from the fully excited state $\rho(0)$ can be described in a linearized picture, in which the two jump operators $J_{\pm}$ act independently. At a later stage, the linearized picture is no longer appropriate, since the states generated by multiple applications of $J_{\pm}$ are not orthogonal.
@@ III.1 Discrete Truncated Wigner Approximation
Figure 2: Numerical methods. (a) Sketch of the extended system described by Eq. (7), where the collective decay into left- and right-going waveguide modes is realized effectively via the coupling of the atoms to two lossy cavity modes with decay rate $\kappa$. This extended model underlies the DTWA method discussed in Sec. () and is used for the study of non-Markovian effects in Sec. VI.2. In the Markovian limit $\kappa/(\gamma N)\to\infty$, the model recovers the original Lindblad master equation in Eq. (7). (b) Sketch of a homodyne measurement setup that corresponds to the QSD unraveling of the master equation, as discussed in Sec. III.2. Trajectories obtained in our QSDMF simulations represent (factorized) states of the atoms that are conditioned on a given history of the measured homodyne currents $I_{\rm R}(t)$ and $I_{\rm L}(t)$.
@@
$\begin{split}\frac{d}{dt}s_{j}^{-}&=i\frac{\gamma}{2}\sum_{l\neq j}\sin(|\xi_{j}-\xi_{l}|)s_{j}^{z}s_{l}^{-}\\ &+i\frac{g}{\sqrt{2}}s_{j}^{z}\left(\alpha_{\text{R}}e^{i\xi_{j}}+\alpha_{\text{L}}e^{-i\xi_{j}}\right),\\ \frac{d}{dt}s_{j}^{z}&=-\gamma\sum_{l\neq j}\sin(|\xi_{j}-\xi_{l}|)(is_{j}^{+}s_{l}^{-}-is_{j}^{-}s_{l}^{+})\\ &-\sqrt{2}gs_{j}^{+}\left(\alpha_{\text{R}}e^{i\xi_{j}}+\alpha_{\text{L}}e^{-i\xi_{j}}\right)-{\text{c.c.}},\end{split}$ (9)
@@ III.2 Quantum State Diffusion
As depicted in Fig. 2(b), QSD describes the dynamics of the monitored waveguide system under homodyne measurement of the emitted field amplitudes. In this formalism, the master equation (1) is unraveled into an ensemble of pure trajectories obeying the QSD equation [53, 54]
@@
with $k=\text{R, L}$ and $dW_{k}$ two complex-valued Wiener increments. Each trajectory represents the state of the system conditioned on the homodyne currents
$dI_{k}(t)=\frac{\gamma}{2}\langle J_{k}+J_{k}^{\dagger}\rangle dt+\sqrt{\frac{\gamma}{2}}dW_{k}$ (16)
@@
In general, the wavefunction in Eq. (III.2) still scales exponentially with the atom number, and to make this method tractable, we approximate each trajectory with a product state, $|\psi\rangle\approx\otimes_{j=1}^{N}|\psi_{j}\rangle$. This simplification allows us to describe each trajectory again in terms of classical spin variables $\vec{\mathfrak{s}}_{j}(t)$, with components $\mathfrak{s}^{\eta}_{j}=\langle\psi_{j}|\sigma_{j}^{\eta}|\psi_{j}\rangle$. The corresponding stochastic equations for these spin vectors for the current setting are given in Ref. [36]. Note that QSDMF can be improved by refining the treatment of individual trajectories. For example, including second-order moments [59, 60] in the EOM of trajectories can encode some quantum correlations, at the cost of a less favorable scaling $\mathcal{O}(N^{2})$. Importantly, by treating trajectories using matrix product states [36], it can be systematically improved by increasing the bond dimension $d$, with complexity $\mathcal{O}(Nd^{3})$. This method is thus particularly efficient if entanglement in each trajectory is constant or logarithmic in $N$, which requires $d\sim\mathcal{O}(1)$ or $d\sim\mathcal{O}(N)$ respectively for giving numerically exact results.
@@ III.4 Comparison with Other Methods
Methods based on a cumulant expansion (“CE”) of correlations [61, 62, 24] scale as $\mathcal{O}(N^{C})$, where $C\geq 2$ is the truncation order. As a result, such simulations become costly for large systems. Both our DTWA approach and CE methods can handle non-Markovian effects, for example, by coupling the atoms to additional cavity modes. However, as shown in Sec. VI.2 and Appendix C, in this non-Markovian setting, CE requires $C>3$ to converge, which increases the computational cost significantly. On the other hand, our DTWA method is still very accurate while keeping $\mathcal{O}(N)$ complexity.
To conclude, out of these various methods, the DTWA approach described above combines an efficient scaling, $\mathcal{O}(N)$, with the ability to account for both classical and—up to a certain degree—quantum correlations and permits a straightforward and numerically inexpensive generalization to non-Markovian settings. This makes this method suitable for studying superradiance in a large variety of different scenarios, and has also been selected as the method of choice for most parts of the numerical analysis. Additionally, we use QSDMF to gain trajectory-level insights, for example, on the spontaneous spin ordering studied in Sec. V.
@@ IV Superradiant Scaling in Disordered Arrays
Based on the numerical methods introduced above, we now perform a systematic investigation of the scaling behavior of superradiance for a varying level of disorder. Of central interest in this study is the scaling of the peak emission rate $\mathcal{R}$ and the burst time $t_{\star}$ with the number of atoms, $N$, and how this scaling is affected by disorder. In addition, we consider second-order photon-photon correlations, which are frequently used as an indicator for the onset of superradiance in the initial dynamics, but are less well understood during the subsequent burst dynamics.
@@ IV.1 Superradiant Scaling
Figure 5: Superradiant scaling in the presence of spatial disorder. (a) Plot of the scaled superradiant rate, $\mathcal{R}_{\star}/(\gamma N^{2})$, as a function of the atom number NN and for a varying disorder strength $\Theta$. The plot demonstrates the asymptotic recovery of the Dicke superradiant scaling with finite-size corrections that are fitted to the analytic prediction in Eq. (18) (dashed lines). Similarly, the numerical data presented in (b) shows the corresponding scaled burst time, $t_{\star}\gamma N/\log(N)$, again converging to the asymptotic Dicke scaling for large $N$. In both plots, the results have been obtained using the DTWA method, averaged over $\mathcal{N}=10^{3}$ trajectories.
@@ IV.2 Finite-Size Scaling
To explain the different scalings toward the asymptotic limit, we return to the ‘diagonalized’ form of the decaying channels given in Eq. (5). Further, we make the simplifying assumption that, for a given disorder realization, the decay dynamics can be qualitatively captured by the maximal eigenvalue of the decay matrix, $\Gamma_{+}=\gamma N(1+|c|)$, which is determined by the overlap parameter in Eq. (6). For phases $\xi_{j}$ that are uniformly distributed in $\Theta/2,\Theta/2)$, the mean value $\mu_{a}$ ($\mu_{b}$) and the variance $\sigma_{a}$ ($\sigma_{b}$) for the quantity $\Re e^{i2\xi_{j}}$ ($\Im e^{i2\xi_{j}}$) are given by
@@
$E\left[\Gamma_{+}\right]=\frac{1}{2}\gamma N\left(1+\sqrt{\frac{\pi}{8}}N^{-1/2}\right),$ (21)
in the limit of large $N$. This gives a different finite-size correction $\sim N^{-1/2}$.
While this analysis does not allow us to predict the numerical prefactor $r_{0}$, it shows that the observed scaling relations for the maximal decay rate during the burst can be obtained from the relation $\mathcal{R}_{\star}\propto N\Gamma_{+}$. This observation is rather surprising, as in view of the nonlinear dynamics or the presence of additional Hamiltonian terms, such a simple correspondence is a priori not expected. As shown in Appendix A and discussed below, this observation is, however, consistent with an upper bound that is proportional to $\Gamma_{+}N$.

### v2

# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide Quantum Electrodynamics

@@ I Introduction
Recently, there has been a revived interest in superradiance [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], mainly stimulated by the development of new experimental platforms, in which superradiant effects can be explored at an unprecedented level of control, and potentially be exploited in applications such as superradiant lasers [43, 44, 45] or quantum metrology [46, 47, 48, 49]. Of particular relevance in this context are waveguide QED systems [50, 51, 52], where the emitted photons are confined to a one-dimensional (1D) channel and atom-atom correlations can be established efficiently and over long distances [16, 18, 19, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Specifically, when emitters are placed in a lattice with a spacing that matches an exact multiple of the resonant wavelength, all the emitted photons can interfere fully constructively, thereby recovering the features of the original Dicke model in an extended setting. However, in real systems, the positions of the atoms as well as their frequencies fluctuate, giving rise to competing interference patterns once the disorder becomes sufficiently strong. This raises an important fundamental question: Does Dicke superradiance survive in disordered quantum systems, and how does disorder affect its characteristic features?
Figure 1: (a) Sketch of a disordered waveguide QED system, where NN excited TLAs are located at random positions $z_{j}$ along a 1D photonic channel and undergo collective spontaneous decay. As shown in the inset, the displacements $\delta_{j}$ from exact multiples of the resonant wavelength $\lambda_{0}$ are chosen randomly from the interval $\Theta/2,\Theta/2)$ to interpolate between a regular array ($\Theta=0)$ and a uniformly distributed gas of atoms ($\Theta\gg 1$). (b) Illustration of the two ordering patterns (red and blue), which are generated spontaneously between the atomic dipoles during the superradiant decay. The dipoles self-align according to their distances, with $\phi_{j}\approx\phi_{1}+k_{0}(z_{j}-z_{1})$ (blue arrows) or $\phi_{j}\approx\phi_{1}-k_{0}(z_{j}-z_{1})$ (red arrows), which even for strong disorder yields constructive interference and enhanced collective emission either to the left or to the right.
@@
By investigating the maximal emission rate and the peak emission time for an increasing number of atoms, we demonstrate that even for strong disorder, the Dicke scaling is asymptotically recovered. There are, however, subtle differences in the numerical prefactors and in how this limit is approached. To explain these observations, we introduce a product-state ansatz for the configuration of atomic dipoles at the time of the emission peak, which provides a tight upper bound for the maximal collective decay rate. Remarkably, this bound is not only close to the numerically exact values, but also reveals a simple and intuitive mechanism for robustness: the atomic spins organize spontaneously in random directions, but with their relative orientations being determined by their relative locations along the waveguide [Fig. 1(b)]. This emergent spin ordering, without any external control, enables constructive interference in the decay channels for any atomic configuration, which enables the persistence of the $N^{2}$ scaling regardless of the disorder strength.
@@ II Model
We consider a set of $N$ TLAs, which are located at fixed positions $z_{j}$ along a 1D photonic waveguide [see Fig. 1(a)]. The TLAs have a ground state $|g\rangle$ and an excited state $|e\rangle$, which are separated by a transition frequency $\omega_{0}$. Under the assumption that over a sufficiently broad frequency range around $\omega_{0}$ the waveguide exhibits an approximately linear dispersion relation with group velocity $v_{\rm g}$, we can trace out the waveguide modes and describe the remaining dynamics of the atoms by a master equation of the form (see, for example, [57, 58])
@@
In our study, the offset phases $\xi_{j}=k_{0}\delta_{j}$ are randomly chosen from a uniform distribution in the interval $\Theta/2,\Theta/2)$, such that $\Theta$ quantifies the degree of disorder. Note that for this configuration, neighboring atoms can overlap when $\Theta>2\pi$ and in the limit $\Theta\rightarrow\infty$ we recover a fully homogeneous density distribution spatially. The dynamics is determined by the distribution of the relative propagation phases, $\Delta\xi_{ij}=(\xi_{i}-\xi_{j})\mod 2\pi$, which obeys a wrapped triangular distribution (see the Supplemental Material [60]) and approaches a uniform distribution for any $\Theta=2n\pi$ with $n\in\mathbb{Z}_{+}$ and in the limit $\Theta\rightarrow\infty$. Additional simulations for an ensemble with Gaussian-distributed $\xi_{j}$ are performed for comparison, ensuring that the general conclusions do not depend on any specifics of the disorder.
@@
where $J_{\pm}$ are collective jump operators and $\Gamma_{\pm}=\gamma N(1\pm|c|)/2$ are the corresponding decay rates. Here, we have introduced the complex parameter
@@
which characterizes the overlap between the original left- and right-going decay modes (see Appendix A). It assumes a value of $c=1$ in the absence of disorder, i.e., $\xi_{j}=0$, and a value of $c\simeq 0$ when the atoms are homogeneously distributed. The short-time dynamics from the fully excited state $\rho(0)$ can be described in a linearized picture, in which the two jump operators $J_{\pm}$ act independently. At a later stage, the linearized picture is no longer appropriate, since the states generated by multiple applications of $J_{\pm}$ are not orthogonal.
@@ III.1 Discrete Truncated Wigner Approximation
Figure 2: Numerical methods. (a) Sketch of the extended system described by Eq. (7), where the collective decay into left- and right-going waveguide modes is realized effectively via the coupling of the atoms to two lossy cavity modes with decay rate $\kappa$. This extended model underlies the DTWA method discussed in Sec. III.1 and is used for the study of non-Markovian effects in Sec. VI.2. In the Markovian limit $\kappa/(\gamma N)\to\infty$, the model recovers the original Lindblad master equation in Eq. (7). (b) Sketch of a heterodyne measurement setup that corresponds to the QSD unraveling of the master equation, as discussed in Sec. III.2. Trajectories obtained in our QSDMF simulations represent (factorized) states of the atoms that are conditioned on a given history of the measured heterodyne currents $I_{\rm R}(t)$ and $I_{\rm L}(t)$.
@@
$\begin{split}\frac{d}{dt}s_{j}^{-}&=i\frac{\gamma}{2}\sum_{l\neq j}\sin(|\xi_{j}-\xi_{l}|)s_{j}^{z}s_{l}^{-}\\ &+\frac{g}{\sqrt{2}}s_{j}^{z}\left(\alpha_{\text{R}}e^{i\xi_{j}}+\alpha_{\text{L}}e^{-i\xi_{j}}\right),\\ \frac{d}{dt}s_{j}^{z}&=-\gamma\sum_{l\neq j}\sin(|\xi_{j}-\xi_{l}|)(is_{j}^{+}s_{l}^{-}-is_{j}^{-}s_{l}^{+})\\ &-\sqrt{2}gs_{j}^{+}\left(\alpha_{\text{R}}e^{i\xi_{j}}+\alpha_{\text{L}}e^{-i\xi_{j}}\right)-{\text{c.c.}},\end{split}$ (9)
@@ III.2 Quantum State Diffusion
As depicted in Fig. 2(b), QSD describes the dynamics of the monitored waveguide system under heterodyne measurement of the emitted field amplitudes. In this formalism, the master equation in Eq. (1) is unraveled into an ensemble of pure trajectories obeying the QSD equation [55, 56]
@@
with $k=\text{R, L}$ and $dW_{k}$ two complex-valued Wiener increments. Each trajectory represents the state of the system conditioned on the heterodyne currents
$dI_{k}(t)=\frac{\gamma}{2}\langle J_{k}\rangle dt+\sqrt{\frac{\gamma}{2}}dW_{k}$ (16)
@@
In general, the wavefunction in Eq. (III.2) still scales exponentially with the atom number, and to make this method tractable, we approximate each trajectory by a product state, $|\psi\rangle\approx\otimes_{j=1}^{N}|\psi_{j}\rangle$. This simplification allows us to describe each trajectory again in terms of classical spin variables $\vec{\mathfrak{s}}_{j}(t)$, with components $\mathfrak{s}^{\eta}_{j}=\langle\psi_{j}|\sigma_{j}^{\eta}|\psi_{j}\rangle$. The corresponding stochastic equations for these spin vectors for the current setting are given in Ref. [36]. Note that QSDMF can be improved by refining the treatment of individual trajectories. For example, including second-order moments [62, 63] in the EOM of trajectories can encode some quantum correlations, at the cost of a less favorable scaling $\mathcal{O}(N^{2})$. Importantly, by treating trajectories using matrix product states [36], it can be systematically improved by increasing the bond dimension $d$, with complexity $\mathcal{O}(Nd^{3})$. This method is thus particularly efficient if entanglement in each trajectory is constant or logarithmic in $N$, which requires $d\sim\mathcal{O}(1)$ or $d\sim\mathcal{O}(N)$, respectively, for giving numerically exact results.
@@ III.4 Comparison with Other Methods
Methods based on a cumulant expansion (CE) of correlations [64, 65, 24] scale as $\mathcal{O}(N^{C})$, where $C\geq 2$ is the truncation order. As a result, such simulations become costly for large systems. Both our DTWA approach and CE methods can handle non-Markovian effects, for example, by coupling the atoms to additional cavity modes. However, as shown in Sec. VI.2 and Appendix C, in this non-Markovian setting, CE requires $C>3$ to converge, which increases the computational cost significantly. On the other hand, our DTWA method is still very accurate while keeping $\mathcal{O}(N)$ complexity.
To conclude, out of these various methods, the DTWA approach described above combines an efficient scaling, $\mathcal{O}(N)$, with the ability to account for both classical and—up to a certain degree—quantum correlations and permits a straightforward and numerically inexpensive generalization to non-Markovian settings. This makes this method suitable for studying superradiance in a large variety of different scenarios, and has also been selected as the method of choice for most parts of the numerical analysis below. Additionally, we use QSDMF to gain trajectory-level insights, for example, on the spontaneous spin ordering studied in Sec. V.
@@ IV Superradiant Scaling in Disordered Arrays
Based on the numerical methods introduced above, we now perform a systematic investigation of the scaling behavior of superradiance for a varying level of disorder and setting $z_{j}=\delta_{j}$  [^2] While displacing the atomic positions by $z_{j}\rightarrow z_{j}+m_{j}\times\lambda/2$ with $m_{j}\in\mathbbm{Z}$ does not affect the dissipators of our model, it can alter some terms of the Hamiltonian. However, this only shifts the precise values of $\mathcal{R}_{\star}/(\gamma N^{2})$ by at most a few percent and leaves all qualitative features unchanged, as shown in both numerical simulations and the analytical analysis later. The results obtained for $z_{j}=\delta_{j}$ are thus representative for all configurations where atoms are spaced by multiples of $\lambda/2$ on average. Of central interest in this study is the scaling of the peak emission rate $\mathcal{R}_{\star}$ and the burst time $t_{\star}$ with the number of atoms, $N$, and how this scaling is affected by disorder. In addition, we consider second-order photon-photon correlations, which are frequently used as an indicator for the onset of superradiance in the initial dynamics, but are less well understood during the subsequent burst dynamics.
@@ IV.1 Superradiant Scaling
Figure 5: Superradiant scaling in the presence of spatial disorder. (a) Plot of the scaled superradiant rate, $\mathcal{R}_{\star}/(\gamma N^{2})$, as a function of the atom number NN and for a varying disorder strength $\Theta$. The plot demonstrates the asymptotic recovery of the Dicke superradiant scaling with finite-size corrections that are fitted to the analytic prediction in Eq. (18) (solid lines). Similarly, the numerical data presented in (b) shows the corresponding scaled burst time, $t_{\star}\gamma N/\log(N)$, which again converges to the asymptotic Dicke scaling for large $N$. In both plots, the results have been obtained using the DTWA method, averaged over $\mathcal{N}=10^{3}$ trajectories.
@@ IV.2 Finite-Size Scaling
To explain the different scalings toward the asymptotic limit, we return to the diagonalized form of the decay channels given in Eq. (5). Further, we make the simplifying assumption that, for a given disorder realization, the decay dynamics can be qualitatively captured by the maximal eigenvalue of the decay matrix, $\Gamma_{+}=\gamma N(1+|c|)$, which is determined by the overlap parameter in Eq. (6). For phases $\xi_{j}$ that are uniformly distributed in $\Theta/2,\Theta/2)$, the mean value $\mu_{a}$ ($\mu_{b}$) and the variance $\sigma_{a}$ ($\sigma_{b}$) for the quantity $\Re e^{i2\xi_{j}}$ ($\Im e^{i2\xi_{j}}$) are given by
@@
$E\left[\Gamma_{+}\right]=\frac{1}{2}\gamma N\left(1+\sqrt{\frac{\pi}{8N}}\right),$ (21)
in the limit of large $N$. This gives a different finite-size correction $\sim N^{-1/2}$. We remark that this situation already occurs for $\Theta=\pi$, where for the situation with regular lattice spacing $\lambda$, the atoms are still spaced by at least $\lambda/2$. The reason is that this remaining gap corresponds to a propagation phase of $k_{0}\lambda/2=\pi$ between neighboring atoms. Then this phase $\pi$ does not change the maximal decay rate $\Gamma_{+}$, since $c$ depends on $2\xi_{j}$.
While this analysis does not allow us to predict the numerical prefactor $r_{0}$, it shows that the observed scaling relations for the maximal decay rate during the burst can be obtained from the relation $\mathcal{R}_{\star}\propto N\Gamma_{+}$. This observation is rather surprising, as in view of the nonlinear dynamics or the presence of additional Hamiltonian terms, such a simple correspondence is a priori not expected. As discussed in Appendix E, this observation is, however, consistent with an upper bound that is proportional to $\Gamma_{+}N$.